OpenStudy (anonymous):

University Theater sold 556 tickets for a play. Tickets cost $22 per adult 2) and $12 per senior citizen. If total receipts were $8492, how many senior citizen tickets were sold? (I have to use X and Y to answer the Question)

OpenStudy (anonymous):

this is Question 2 can anyone please help me with the 2 problems please... A flat rectangular piece of aluminum has a perimeter of 60 inches. The 9) length is 14 inches longer than the width. Find the width.

OpenStudy (anonymous):

First you know that 556=x+y and 22x+12y=8492

OpenStudy (anonymous):

Now you need to get your initial function into one term so that you can substitute it into the second equation

OpenStudy (anonymous):

556-x=y then sub for y 22x+12(556-x)=8492 solve for x and that will give you the number of adult tickets then with that x plug it back into the 556-x=y and the y that you get is the number of senior tickets